# StrandHead: Text to Hair-Disentangled 3D Head Avatars Using Human-Centric Priors

### v3

@@ 2 Related Work
Task Method
Head-Hair-
Decoupled
Strand-Based
Hair
Geometry &
Texture
No Large-Scale
3D-Text Paired Data
Text-to-Head
[12, 30, 94, 17]
[84]
Text-to-Hair
[62, 80, 28, 48]
[65]
Text-to-Head-Hair
Ours
@@ 3.1 Preliminaries
FLAME [26] is a 3D parametric head model that represents the human head shape, pose, and expression using a compact set of parameters. Given a set of shape parameters $\beta$, pose parameters ${\theta}_{\text{pose}}$, and expression parameters $\psi_{\text{exp}}$, FLAME can create a 3D head mesh ${M}$.
@@ 3.4 Hair Geometry Generation
[…]
@@
$\mathcal{L}_{\text{hair-geo}}=$ $\mathcal{L}^{\text{hd}}_{SDS}+\lambda^{\text{hn}}_{SDS}\mathcal{L}^{\text{hn}}_{SDS}+\lambda_{\text{ori}}\mathcal{L}_{\text{ori}}+\lambda_{\text{cur}}\mathcal{L}_{\text{cur}}$
@@ 6.2 Implementation Details
Hyper-Parameters. For the bald head generation, our approach requires 10,000 iterations for geometry creation, 2,000 iterations for texture synthesis, and 5,000 iterations for appearance refinement using MSDS [17]. The loss weight $\lambda_{\text{prior}}$ is set to $1\times 10^{3}$. For hair generation, the process involves 5,000 iterations for geometry modeling, 2,000 iterations for texture generation, and 5,000 iterations for visual enhancement using MSDS [17]. The respective loss weights for $\lambda^{\text{hn}}_{SDS}$, $\lambda_{\text{ori}}$, $\lambda_{\text{cur}}$, $\lambda_{\text{bbox}}$, $\lambda_{\text{face}}$, and $\lambda_{\text{colli}}$ are assigned as $1\times 10^{-3}$, $1\times 10^{4}$, $1\times 10^{4}$, $1\times 10^{3}$, $1\times 10^{3}$, and $1\times 10^{3}$ respectively. The number of strand polylines $N_{s}$ and strand points $N_{p}$ are configured as 3000 and 100, respectively. For different hairstyles—normal, straight, wavy, and curly—the target curvature $C_{\text{target}}$ is defined as $5\times 10^{-2}$, $2\times 10^{-2}$, $1\times 10^{-1}$, and $2\times 10^{-1}$, respectively. All experiments are conducted on an Ubuntu server equipped with A6000 GPUs. Generating a bald head takes roughly 4 hours with 24GB of memory, and generating a haircut takes roughly 4 hours with 44GB of memory.
@@
where $\mathcal{L}_{\text{chamf}}$ is the Chamfer distance between $\hat{M}_{\text{init}}$ and the current human head geometry, $\mathcal{L}_{\text{edge}}$ is the edge length regularization loss, $\mathcal{L}_{\text{nor}}$ is the normal consistency loss, and $\mathcal{L}_{\text{lap}}$ is the Laplacian smoothness loss. The loss weights $\lambda_{\text{chamf}}$, $\lambda_{\text{edge}}$, $\lambda_{\text{nor}}$, and $\lambda_{\text{lap}}$ are set to $1\times 10^{2}$, $1\times 10^{0}$, $1\times 10^{-2}$, and $1\times 10^{-1}$, respectively. The ablation study results for the FLAME-evolving prior loss are presented in Fig. 12. As shown, excluding $\mathcal{L}_{\text{prior}}$ results in exaggerated head shapes (e.g., overly pointed head tops), significantly compromising the realism of the outputs. Employing a non-evolving prior loss achieves reasonable head proportions but fails to preserve finer geometric details such as beards and wrinkles. By leveraging the full $\mathcal{L}_{\text{prior}}$, our approach generates head geometry with both accurate proportions and high detail fidelity.

### v4

@@ 2 Related Work
Task Method Head-Hair- Decoupled Strand-Based Hair Geometry & Texture No Large-Scale 3D-Text Paired Data
Text-to-Head [12, 30, 94, 17] [84]
Text-to-Hair [62, 80, 28, 48] [65]
Text-to-Head-Hair Ours
@@ 3.1 Preliminaries
FLAME [26] is a 3D parametric head model that represents the human head shape, pose, and expression using a compact set of parameters. Given a set of shape parameters $\beta$, pose parameters ${\theta}_{\text{pose}}$, and expression parameters $\psi_{\text{exp}}$, FLAME can create a 3D head mesh M{M}.
@@ 3.4 Hair Geometry Generation
where $A(i)$ is the set of adjacent strands for the $i$-th strand, and $|A(i)|$ is the number of neighboring strands.
@@
$\mathcal{L}_{\text{hair-geo}}=\$$\mathcal{L}^{\text{hd}}_{SDS}+\lambda^{\text{hn}}_{SDS}\mathcal{L}^{\text{hn}}_{SDS}+\lambda_{\text{ori}}\mathcal{L}_{\text{ori}}+\lambda_{\text{cur}}\mathcal{L}_{\text{cur}}$
@@ 6.2 Implementation Details
Hyper-Parameters. For the bald head generation, our approach requires 10,000 iterations for geometry creation, 2,000 iterations for texture synthesis, and 5,000 iterations for appearance refinement using MSDS [17]. The loss weight $\lambda_{\text{prior}}$ is set to $1\times 10^{3}$. For hair generation, the process involves 5,000 iterations for geometry modeling, 2,000 iterations for texture generation, and 5,000 iterations for visual enhancement using MSDS [17]. The respective loss weights for $\lambda^{\text{hn}}_{SDS}$, $\lambda_{\text{ori}}$, $\lambda_{\text{cur}}$, $\lambda_{\text{bbox}}$, $\lambda_{\text{face}}$, and $\lambda_{\text{colli}}$ are assigned as $\times 10^{-3}, 1×1$$\times 10^{4}$, $\times 10^{4}, 1×1$$\times 10^{3}$, $1\times 10^{3}$, and $1\times 10^{3}$ respectively. The number of strand polylines $N_{s}$ and strand points $N_{p}$ are configured as 3000 and 100, respectively. For different hairstyles—normal, straight, wavy, and curly—the target curvature $C_{\text{target}}$ is defined as $5\times 10^{-2}$, $2\times 10^{-2}$, $1\times 10^{-1}$, and $2\times 10^{-1}$, respectively. All experiments are conducted on an Ubuntu server equipped with A6000 GPUs. Generating a bald head takes roughly 4 hours with 24GB of memory, and generating a haircut takes roughly 4 hours with 44GB of memory.
@@
where $\mathcal{L}_{\text{chamf}}$ is the Chamfer distance between $\hat{M}_{\text{init}}$ and the current human head geometry, $\mathcal{L}_{\text{edge}}$ is the edge length regularization loss, $\mathcal{L}_{\text{nor}}$ is the normal consistency loss, and $\mathcal{L}_{\text{lap}}$ is the Laplacian smoothness loss. The loss weights $\lambda_{\text{chamf}}$, $\lambda_{\text{edge}}$, $\lambda_{\text{nor}}$, and $\lambda_{\text{lap}}$ are set to $\times 10^{2}, 1×1$$\times 10^{0}$, $1\times 10^{-2}$, and $1\times 10^{-1}$, respectively. The ablation study results for the FLAME-evolving prior loss are presented in Fig. 12. As shown, excluding $\mathcal{L}_{\text{prior}}$ results in exaggerated head shapes (e.g., overly pointed head tops), significantly compromising the realism of the outputs. Employing a non-evolving prior loss achieves reasonable head proportions but fails to preserve finer geometric details such as beards and wrinkles. By leveraging the full $\mathcal{L}_{\text{prior}}$, our approach generates head geometry with both accurate proportions and high detail fidelity.